# CDKFormer: Contextual Deviation Knowledge-Based Transformer for Long-Tail Trajectory Prediction

### v1

@@ IV-D2 Dual Query-Based Trajectory Decoder
Then, we combine the dual queries through a learnable gating mechanism. The gate weights are learned through an MLP-based router module. The input of the router module is the concatenated dual queries. The tail weight $\bm{\gamma}\in\mathbb{R}^{T_{f}\times d}$ is generated and subsequently used for dual query combination as follows.
@@
To allow the scene query to fully learn the temporal dependency among difference modalities, we apply self-attention block on time and modality dimension subsequently. The scene query is converted to a dense feature in shape $\mathbb{R}^{KT_{f}\times d}$ before the refinement, similar to [36]. Then, we update the scene query $\bm{Q}$ with another multistream decoder block.
@@
Afterwards, another self-attention based refinement is conducted, followed by an MLP to generate future multimodal trajectories and the corresponding probability scores.
IV-E Learning Objectives
In this study, we employ distinct loss functions to regulate the learning of different query types. Specifically, an smoothed L1 regression loss is applied to the predictions of mode queries, future queries, and scene queries. For mode learning in both mode and scene queries, we utilize a cross-entropy classification loss. The winner-takes-all strategy is adopted. Additionally, we apply a regression loss on the predicted future positions of neighboring agents learned from the contextual feature.
$\mathcal{L}_{\rm mode}$ $=\mathcal{L}_{\rm mode}^{\rm reg}+\mathcal{L}_{\rm mode}^{\rm cls}$ (19)
$\mathcal{L}_{\rm future}$ $=\mathcal{L}_{\rm future}^{\rm reg}$ (20)
$\mathcal{L}_{\rm scene}$ $=\mathcal{L}_{\rm scene}^{\rm reg}+\mathcal{L}_{\rm scene}^{\rm cls}$ (21)
$\mathcal{L}_{\rm grp}$ $=\mathcal{L}_{\rm grp}^{\rm reg}$ (22)
@@ IV-E Learning Objectives
$\mathcal{L}^{\rm reg}_{*}$ =$=\left\{\begin{array}[]{lr}\frac{1}{t_{f}}\sum_{t}^{t+t_{f}}\frac{1}{2}(\hat{y}^{t}-y^{t})^{2},&\text{if}\ |\hat{y}^{t}-y^{t}|<1\\ \frac{1}{t_{f}}\sum_{t}^{t+t_{f}}(|\hat{y}^{t}-y^{t}|-0.5),&\text{o.w.}\end{array}\right.$ (25)
$\mathcal{L}^{\rm cls}_{*}$ $=-\sum_{k=1}^{K}p_{k}\log(\hat{p}_{k})$ (26)
where $\hat{y}^{t}$ and $y^{t}$ is the predicted and ground-truth future position at time $t$, and $p_{k}$ is the probability for selecting mode $k$ as the best mode. $\hat{p}_{k}$ is the ground-truth best prediction index.
The tail query $Q_{t}$ is designed especially to generate accurate future 2D position on tailed samples. Following the idea of label distribution smoothing (LDS) in [44], we introduce a tail loss weighted by smoothed tail score $S$.
$\mathcal{L}_{\rm tail}=\tilde{S}\cdot\mathcal{L}_{\rm tail}^{\rm reg}$ (27)
where $\tilde{S}$ is the tail score smoothed with a Gaussian kernel. $\mathcal{L}_{\rm tail}^{\rm reg}$ is the regression loss on tail query predictions. The hard samples are therefore given higher weights to support imbalanced regression. The final loss function is calculated as follows.
$\mathcal{L}=\mathcal{L}_{\rm mode}+\mathcal{L}_{\rm future}+\mathcal{L}_{\rm scene}+\mathcal{L}_{\rm grp}+\alpha\mathcal{L}_{\rm tail}$ (28)
@@
V Experiments
V-A Experimental Settings
V-A1 Dataset
We use Argoverse 2 motion forecasting dataset [6] to validate the performance of CDKFormer. The Argoverse 2 motion forecasting dataset consists of 250,000 scenarios for training and validation. Each scenario lasts 11 seconds and includes the 2D bird’s-eye view of the centroid and heading of each tracked object, sampled at 10 Hz.
V-A2 Baseline Comparison
@@
We compare our model with the following SOTA trajectory prediction models: VectorNet [26], LaneGCN [25], MTR [20] and QCNet [21].
@@
We also implement two SOTA long-tail trajectory prediction methods as comparison baselines to varify the long-tail performance of the proposed model: Contrastive [13] and FEND [14]. Additionally, we compare two naive long-tail learning approaches: data balanced sampling and loss reweighting. The sampling and reweighting factors are set as the tail scores of each sample. Long-tail learning methods are implemented using QCNet as the backbone.
V-A3 Metrics
We adopt the commonly used trajectory prediction metric: minimum average displacement error (minADEk) and minimum final displacement error (minFDEk) between $K$ trajectory samples and the ground truth to measure the accuracy of trajectory prediction, as calculated below:
$\rm{minADE_{k}}$ $=\frac{1}{t_{f}}\min_{k}\sum_{t=1}^{t_{f}}\|\hat{y}^{t}_{k}-y^{t}\|^{2}$ (29)
$\rm{minFDE_{k}}$ $=\min_{k}\|\hat{y}^{t_{f}}_{k}-y^{t_{f}}\|^{2}$ (30)
where $\hat{y}^{t}_{k}$ denotes the $k$-th predicted future position of the target vehicle at time $t$, and ${y}^{t}$ denotes the ground truth future position at time $t$.
@@ V-A3 Metrics
We further evaluate our model’s performance on both head and tail samples to validate its effectiveness in long-tailed scenarios. Specifically, we report the minADEk and minFDEk scores for the top 5% and 10% tailed samples, which are selected based on the long-tail score $S$.
V-A4 Implementation Details
The model is trained on a server with 4 NVIDIA GeForce RTX 4090 GPUs. We set the embedding dimension to 64. The tail weight $\alpha$ is 0.1. The number of encoder and decoder layers is set to 2. The training is conducted for 30 epochs. The Adam optimizer is used with an initial learning rate of 0.003 and a weight decay rate of 0.01. A coseine learning rate scheduler is employed to adjust the learning rate during training. The batch size is set to 16. The implementation is based on Python 3.9 and PyTorch 2.0.
TABLE II: Prediction performance in comparison with SOTA trajectory prediction models. The results are presented in minADE6/minFDE6.
All Top 10% Top 5%
VectorNet [26] 1.11/2.90 2.06/3.61 2.65/4.60
LaneGCN [25] 1.07/1.84 1.99/3.50 2.55/4.47
MTR [20] 0.86/1.57 1.64/3.05 2.00/3.89
QCNet [21] 0.78/1.41 1.45/2.77 1.77/3.55
CDKFormer 0.75/1.42 1.33/2.58 1.44/2.92
V-B Overall Performance
@@
We fist compare the performance of our model with SOTA models on Argoverse 2 motion forecasting dataset (validation set). The results are presented in Table II.
@@
CDKFormer achieves a minADE6 of 0.75 m and a minFDE6 of 1.42 m on all samples, achieving comparable performance to SOTA models. Compared to a recent strong baseline QCNet [21], the proposed model achieves a 2.56% reduction in minADE6. And a 9.55% reduction in minFDE6 compared to MTR [20] is observed. Compared to earlier graph-based models, our model demonstrates significant performance improvements. For example, when compared to LaneGCN [25], CDKFormer achieves 22.83% lower minFDE6, which can be attributed to the design of our attention-based feature fusion technique and query-based decoding paradigm.
V-C Long-Tail Performance
@@
TABLE III: Prediction performance in comparison with SOTA long-tail learning methods. The results are presented in minADE6/minFDE6.
All Top 10% Top 5%
QCNet [21] 0.78/1.41 1.45/2.77 1.77/3.55
QCNet + Balanced Sampling 0.80/1.45 1.45/2.74 1.74/3.48
QCNet + Loss reweighting 0.79/1.44 1.45/2.75 1.76/3.50
QCNet + Contrastive [13] 0.78/1.43 1.44/2.74 1.75/3.49
QCNet + FEND [14] 0.77/1.43 1.43/2.73 1.73/3.46
CDKFormer 0.75/1.42 1.33/2.58 1.44/2.92
@@
The long-tail performance of CDKFormer is promising compared to SOTA trajectory prediction models, including long-tail prediction methods. As shown in Table II, our model demonstrates significant improvements on the top 10% and top 5% tailed samples, achieving a minFDE6 of 2.58 m and 2.92 m, respectively, outperforming QCNet and other recent models in both metrics by a large margin.
@@
Furthermore, our model exhibits considerable improvements over existing SOTA long-tail learning methods. As illustrated in Table III, traditional imbalanced learning methods show suboptimal performance. Specifically, balanced sampling results in a minADE6 of 1.74 m on topsamples, while loss reweighting achieves a minADE6 of 1.45 m on top 5% samples. Contrastive learning-based methods, such as Contrastive [13] and FEND [14], exhibit limited improvements in long-tail performance. For instance, contrastive learning achieves a minADE6 of 1.44 m and a minFDE6 of 2.74 m on top 10% samples. For the top 5% tailed samples, the proposed CDKFormer achieves a 15.61% improvement on minFDE6, significantly surpassing FEND by a large margin.
V-D Ablation Study
V-D1 Ablation Study on Input Deviation Feature

### v2

@@ 4.4.2 Dual Query-Based Trajectory Decoder
Then, we combine the dual queries through a learnable gating mechanism. Gate weights are learned through an MLP-based router module. The input of the router module is the concatenated dual queries. The tail weight $\bm{\gamma}\in\mathbb{R}^{T_{f}\times d}$ is generated and subsequently used for the dual query combination as follows.
@@
To allow the scene query to fully learn the temporal dependency among difference modalities, we apply self-attention block on time and modality dimension subsequently. The scene query is converted to a dense feature in shape $\mathbb{R}^{KT_{f}\times d}$ before refinement, similar to [zhang2024decoupling]. Then, we update the scene query $\bm{Q}$ with another multistream decoder block.
@@
Afterward, another self-attention-based refinement is performed, followed by an MLP to generate future multimodal trajectories and the corresponding probability scores. Additionally, to ensure each query type learns its representation effectively, auxiliary prediction heads are also attached to the intermediate queries, including the mode query, regular future query and tail future query, to generate future trajectory predictions as well. The detailed loss calculation is provided in the following subsection.
4.5 Learning Objectives
Figure 9: Supervision mechanism for the multicomponent loss. Our model employs deep supervision, where intermediate trajectory predictions are generated from the mode, regular future, tail future queries and the context feature. These outputs are used to compute their respective loss terms ($\mathcal{L}_{\text{mode}}$, $\mathcal{L}_{\text{future,r}}$, $\mathcal{L}_{\text{future,t}}$ and $\mathcal{L}_{\text{group}}$). And the final output are used to compute the scene loss $\mathcal{L}_{\text{scene}}$. The prediction heads are omitted from the diagram for simplicity.
As illustrated in Figure 9, our model is trained using a multicomponent loss function, where each component is designed to supervise a distinct intermediate or final output from our query-based decoder. Specifically, four loss terms regulate the learning of the different query types: $\mathcal{L}_{\text{mode}}$, $\mathcal{L}_{\text{future,r}}$, $\mathcal{L}_{\text{future,t}}$, and $\mathcal{L}_{\text{scene}}$. Due to their multimodal nature, the mode query and the scene query are supervised by composite losses that include both a regression component for trajectory accuracy and a classification component for mode selection. The regular and tail future queries are supervised by regression-only terms.
Crucially, to enhance performance on challenging tail samples, the loss for tail future query $\mathcal{L}_{\text{future,t}}$ is weighted by the smoothed tail score $\tilde{S}$. $\tilde{S}$ is generated from the original tail score $S$ with a Gaussian kernel, following the idea of label distribution smoothing [yang2021delving]. The hard samples are therefore given higher weights to support imbalanced regression. Additionally, an auxiliary regression loss, $\mathcal{L}_{\text{group}}$, is applied to the predicted trajectories of neighboring agents, which are decoded from the first $N_{a}$ dimension of context feature $C_{\rm ctx}$, to encourage robust contextual feature learning.
$\mathcal{L}_{\text{mode}}$ $=\mathcal{L}_{\text{mode}}^{\text{reg}}+\mathcal{L}_{\text{mode}}^{\text{cls}}$ (18)
$\mathcal{L}_{\text{future,r}}$ $=\mathcal{L}_{\text{future,r}}^{\text{reg}}$ (19)
$\mathcal{L}_{\text{future,t}}$ $=\tilde{S}\cdot\mathcal{L}_{\text{future,t}}^{\text{reg}}$ (20)
$\mathcal{L}_{\text{scene}}$ $=\mathcal{L}_{\text{scene}}^{\text{reg}}+\mathcal{L}_{\text{scene}}^{\text{cls}}$ (21)
$\mathcal{L}_{\text{group}}$ $=\mathcal{L}_{\text{group}}^{\text{reg}}$ (22)
For all regression components ($\mathcal{L}_{*}^{\text{reg}}$), we use a smoothed L1 loss. For all classification components ($\mathcal{L}_{*}^{\text{cls}}$), we use a cross-entropy loss. The winner-takes-all strategy is adopted for multimodal predictions.
@@ 4.5 Learning Objectives
$\mathcal{L}^{\text{reg}}_{*}$ $=\left\{\begin{array}[]{ll}\frac{1}{T_{f}}\sum_{t=1}^{T_{f}}\frac{1}{2}({\color[rgb]{0,0,0}\hat{y_{*}}^{t}}-{\color[rgb]{0,0,0}y_{*}^{t}})^{2},&\text{if }|{\color[rgb]{0,0,0}\hat{y_{*}}^{t}}-{\color[rgb]{0,0,0}y_{*}^{t}}|<1\\ \frac{1}{T_{f}}\sum_{t=1}^{T_{f}}(|{\color[rgb]{0,0,0}\hat{y_{*}}^{t}}-{\color[rgb]{0,0,0}y_{*}^{t}}|-0.5),&\text{otherwise}\end{array}\right$. (25)
$\mathcal{L}^{\text{cls}}_{*}$ $=-\sum_{k=1}^{K}{\color[rgb]{0,0,0}p_{*,k}}\log({\color[rgb]{0,0,0}\hat{p}_{*,k}})$ (26)
where $\hat{y}_{*}^{t}$ and $y_{*}^{t}$ are the predicted and ground-truth future positions at time. For the classification loss, $\hat{p}_{*,k}$ and $p_{*,k}$ are the predicted and ground-truth probability label for mode $k$. The subscript * denotes the specific type of output being supervised. The final loss function is calculated as follows.
$\mathcal{L}=\mathcal{L}_{\rm mode}+\mathcal{L}_{\rm future,r}+\alpha\mathcal{L}_{\rm future,t}+\mathcal{L}_{\rm scene}+\mathcal{L}_{\rm group}$ (27)
@@
5 Experiments
5.1 Experimental Settings
5.1.1 Datasets
We use Argoverse 2 motion forecasting dataset [wilson2023argoverse] and inD dataset [bock2020ind] to validate the performance of CDKFormer. The Argoverse 2 motion forecasting dataset consists of 250,000 scenarios, and is officially split into a training set of 200,000 scenarios, a validation set of 25,000 scenarios, and a test set of 25,000 scenarios. Each scenario provides 11 seconds of track histories, with the first 5 seconds as observation and the next 6 seconds as the ground truth for prediction. The data for each tracked object includes its 2D position, heading, velocity, and object type, all sampled at 10 Hz. Vectorized HD map data, including lane boundaries, traffic direction, and crosswalks, are also provided for each scenario. The inD dataset provides trajectories including 2D position, velocity, and heading for various agent types (vehicles, two-wheelers, pedestrians, etc.) captured by drones, sampled at 25 Hz. The dataset consists of 10 hours of measurement data from four intersections in Germany. For our experiments, scenarios were extracted from the continuous recordings using a sliding window approach. We split the dataset into 24,276 training scenarios and 2,547 validation scenarios. Only vehicles were selected as the focal agent for prediction, and both the historical observation window and the prediction horizon were set to 4 seconds.
5.1.2 Baseline Comparison
@@
We compare our model with the following SOTA trajectory prediction models: VectorNet [gao2020vectornet], LaneGCN [liang2020learning], MTR [shi2022motion], QCNet [zhou2023query], HPNet [tang2024hpnet] and LAFormer [liu2024laformer].
@@
We also implement two SOTA long-tail trajectory prediction methods as comparison baselines to varify the long-tail performance of the proposed model: Contrastive [makansi2021exposing] and FEND [wang2023fend]. Additionally, we compare two long-tail learning approaches: data-balanced sampling and loss reweighting. The sampling and reweighting factors are set as the tail scores of each sample. Long-tail learning methods are implemented using QCNet as the backbone.
5.1.3 Metrics
We measure the accuracy of trajectory prediction using several commonly adopted metrics, including minimum average displacement error (minADEk), minimum final displacement error (minFDEk), brier minimum final displacement error ($\text{b-minFDE}_{k}$) and miss rate ($\rm MR_{k}$).
$\mathrm{minADE_{k}}$ $=\frac{1}{T_{f}}\min_{k}\sum_{t=1}^{T_{f}}\|\hat{y}^{t}_{k}-y^{t}\|^{2}$ (28)
$\mathrm{minFDE_{k}}$ $=\min_{k}\|\hat{y}^{T_{f}}_{k}-y^{T_{f}}\|^{2}$ (29)
${\color[rgb]{0,0,0}\text{b-minFDE}_{k}}$ ${\color[rgb]{0,0,0}=\text{minFDE}_{k}+\frac{1}{K}\sum_{k=1}^{K}(\hat{p}_{k}-p_{k})^{2}}$ (30)
where $\hat{{y}}^{t}_{k}$ is the $k$-th predicted future position at time $t$, ${y}^{t}$ is the ground-truth position at time $t$, $\hat{p}_{k}$ is the predicted confidence for the $k$-th trajectory, and $p_{k}$ is the ground-truth label. MRk is the percentage of scenarios where the model fails to produce a single trajectory among $K$ candidates with a final displacement error below a certain threshold, which is 2.0 m in this study. We evaluate all metrics for $K=6$ modes.
@@ 5.1.3 Metrics
We further evaluate our model’s performance on both head and tail samples to validate its effectiveness in long-tailed scenarios. Specifically, we report the scores for the top 5% and 10% tail samples, which are selected based on the long-tail score $S$.
5.1.4 Implementation Details
Our model is implemented in Python 3.9 and PyTorch 2.0, and trained on a server equipped with 4 NVIDIA GeForce RTX 4090 GPUs. The hidden dimension size for all model components is set to 64. For the encoder and decoder, the number of stacked layers is set to 2. The MultiStream Decoder Block utilizes a MoE layer with $K_{e}=8$ experts. In the final loss function, the tail loss weight $\alpha$ is set to 0.1, determined by a grid search.
The model is trained for a total of 30 epochs using the Adam optimizer, with an initial learning rate of $3\times 10^{-3}$ and a weight decay of 0.01. A cosine learning rate scheduler is employed to adjust the learning rate during training. The training is performed with a batch size of 16. For all experiments, we predict $K=6$ multimodal trajectories.
5.2 Overall Performance
@@
We first compare the performance of our model with SOTA models on Argoverse 2 motion forecasting dataset and inD dataset. The results are presented in Table 2 and Table 3.
Table 2: Prediction performance in comparison with SOTA trajectory prediction models on Argoverse 2 motion forecast dataset. Performance is reported for all samples, top 10% tail samples, and top 5% tail samples.
minADE6 minFDE6 b-minFDE6 MR6
VectorNet [gao2020vectornet] 1.12/2.00/2.53 1.99/3.57/4.36 2.65/4.24/5.06 0.24/0.57/0.72
LaneGCN [liang2020learning] 1.06/1.97/2.50 1.84/3.49/4.19 2.46/4.13/4.84 0.23/0.54/0.68
MTR [shi2022motion] 0.86/1.60/1.93 1.56/2.98/3.74 2.14/3.59/4.37 0.21/0.51/0.65
QCNet [zhou2023query] 0.78/1.45/1.77 1.41/2.77/3.55 2.02/3.38/4.18 0.20/0.48/0.60
HPNet [tang2024hpnet] 0.76/1.41/1.70 1.41/2.75/3.50 2.03/3.37/4.17 0.19/0.47/0.57
LAFormer [liu2024laformer] 0.75/1.40/1.67 1.42/2.65/2.98 2.02/3.30/4.09 0.18/0.45/0.55
CDKFormer 0.75/1.33/1.44 1.42/2.58/2.92 2.06/3.25/3.53 0.19/0.46/0.51
Table 3: Prediction performance in comparison with SOTA trajectory prediction models on inD Dataset. Performance is reported for all samples, top 10% tail samples, and top 5% tail samples.
minADE6 minFDE6 b-minFDE6 MR6
VectorNet [gao2020vectornet] 0.30/1.15/1.23 0.82/3.26/3.51 1.20/3.95/4.21 0.13/0.60/0.64
LaneGCN [liang2020learning] 0.31/1.20/1.27 0.83/3.36/3.59 1.16/4.06/4.29 0.14/0.62/0.66
MTR [shi2022motion] 0.27/1.13/1.22 0.73/3.17/3.50 1.04/3.73/4.10 0.12/0.61/0.70
QCNet [zhou2023query] 0.24/1.08/1.19 0.69/3.12/3.44 0.98/3.58/3.95 0.11/0.60/0.69
HPNet [tang2024hpnet] 0.25/1.07/1.16 0.70/3.09/3.37 0.93/3.59/3.98 0.12/0.62/0.69
LAFormer [liu2024laformer] 0.23/0.96/1.04 0.64/2.76/3.01 0.86/3.46/3.72 0.10/0.49/0.55
CDKFormer 0.26/0.92/0.97 0.62/2.44/2.59 0.87/3.14/3.29 0.10/0.45/0.48
@@
As shown in Table 2, our model ties with the recent LAFormer [liu2024laformer] for the best minADE6 at 0.75 m. Compared to another strong baseline QCNet [zhou2023query], the proposed model achieves a 3.85% reduction in minADE6. And an 8.97% reduction in minFDE6 is observed compared to MTR [shi2022motion]. Compared to earlier graph-based models, our model shows significant performance improvements. For example, CDKFormer achieves a b-minFDE6 of 2.06, 16.26% lower than LaneGCN [liang2020learning]. This can be attributed to the design of our attention-based feature fusion technique and query-based decoding paradigm. This SOTA performance is also confirmed on the inD dataset. CDKFormer achieves the best minFDE6 of 0.62 m among all compared methods. Furthermore, it ties with LAFormer for the best MR6 at 0.10. These results validate that CDKFormer performs on par with the leading trajectory prediction models across different datasets and various metrics.
Figure 10: Performance comparison of QCNet and CDKFormer based on minADE6 across various deviation features on Argoverse 2 motion forecast dataset. minADE6 data is categorized into five bins of equal size using quantile binning based on feature values, with results presented as mean values and standard deviation for samples within each bin.
To analyze model performance across the spectrum of each deviation feature, we segmented the validation set into five bins of equal size using quantile binning. Figure 10 plots the mean minADE6 for both CDKFormer and QCNet against the median feature value for each bin, with shaded regions indicating the standard error of the mean. The analysis reveals two key findings. First, for most features, such as $\Delta V_{\text{ind}}$ and $\sigma(H_{\text{grp}})$, the minADE6 for both models tends to increase in the higher-quantile bins, confirming that these features effectively capture the increasing scenario rarity and difficulty. Second, CDKFormer consistently demonstrates a lower mean minADE6 than the QCNet baseline across nearly all features and their respective bins. For instance, in the highest quintile for velocity change ($\Delta V_{\text{ind}}$), CDKFormer achieves a mean minADE6 of 0.79, a 4.82% decrease over QCNet’s 0.83. This trend holds for interaction features as well. For the highest quintile of group heading variance ($\sigma(H_{\text{grp}})$), CDKFormer’s minADE6 of 0.72 is 5.26% lower than QCNet’s 0.76. These findings underscore the use of deviation features in enhancing trajectory prediction accuracy.
5.3 Long-Tail Performance
@@
Table 4: Prediction performance in comparison with SOTA long-tail learning methods on Argoverse 2 motion forecast dataset. Performance is reported for all samples, top 10% tail samples, and top 5% tail samples.
minADE6 minFDE6 b-minFDE6 MR6
QCNet [zhou2023query] 0.78/1.45/1.77 1.41/2.77/3.55 2.02/3.38/4.18 0.20/0.48/0.60
QCNet + Balanced Sampling 0.80/1.43/1.70 1.44/2.72/3.44 2.05/3.29/4.03 0.20/0.47/0.59
QCNet + Loss Reweighting 0.86/1.50/1.78 1.58/2.80/3.43 2.09/3.37/4.12 0.25/0.56/0.65
QCNet + Contrastive [makansi2021exposing] 0.78/1.44/1.75 1.43/2.74/3.49 2.04/3.31/4.02 0.21/0.50/0.62
QCNet + FEND [wang2023fend] 0.77/1.42/1.70 1.42/2.71/3.42 2.03/3.29/3.99 0.20/0.46/0.55
CDKFormer 0.75/1.33/1.44 1.42/2.58/2.92 2.06/3.25/3.53 0.19/0.46/0.51
Table 5: Prediction performance in comparison with SOTA long-tail learning methods on inD Dataset. Performance is reported for all samples, top 10% tail samples, and top 5% tail samples.
minADE6 minFDE6 b-minFDE6 MR6
QCNet [zhou2023query] 0.24/1.08/1.19 0.69/3.12/3.44 0.98/3.58/3.95 0.11/0.60/0.69
QCNet + Balanced Sampling 0.24/1.02/1.10 0.70/2.93/3.17 1.17/3.43/3.67 0.12/0.56/0.61
QCNet + Loss Reweighting 0.29/1.13/1.20 0.82/3.14/3.37 1.07/3.57/3.81 0.17/0.70/0.76
QCNet + Contrastive [makansi2021exposing] 0.24/1.03/1.11 0.68/2.93/3.19 0.91/3.43/3.69 0.11/0.53/0.58
QCNet + FEND [wang2023fend] 0.24/1.02/1.10 0.67/2.89/3.14 0.91/3.40/3.62 0.11/0.52/0.57
CDKFormer 0.26/0.92/0.97 0.62/2.44/2.59 0.87/3.14/3.29 0.10/0.45/0.48
@@
The long-tail performance of CDKFormer is promising compared to SOTA trajectory prediction models, including long-tail prediction methods. As shown in Table 2, on Argoverse 2 motion forecast dataset, our model demonstrates significant improvements on the top 10% and top 5% tail samples, achieving a minFDE6 of 2.58 m and 2.92 m, respectively, outperforming QCNet and other recent models in both metrics by a large margin. This trend also holds true on inD dataset (Figure 3). On top 10% and 5% tail samples, CDKFormer achieved a b-minFDE6 of 3.24 and 3.29, respectively, representing a 12.53% and 17.34% improvement over HPNet [tang2024hpnet].
Figure 11 visualizes the comparison between CDKFormer and QCNet across finely-grained tail score bins on Argoverse 2 motion forecast dataset. Firstly, the minFDE6 values for both models increases as the tail score increases, confirming that the proposed tail score effectively identifies challenging scenarios. Secondly, the performance gap between CDKFormer and QCNet widens significantly in the higher tail score percentiles. While CDKFormer maintains a consistent advantage across all tail bins, its superiority is most pronounced in the most extreme cases. For instance, the relative improvement exceeds 10% for samples in the 91-93% and 96-97% tail score bins and peaks at 17.56% for the most challenging 99-100% bin.
@@
Furthermore, our model exhibits considerable improvements over existing SOTA long-tail learning methods. As illustrated in Table 4 and Table 5, traditional imbalanced learning techniques like loss reweighting and balanced sampling often struggle with a performance trade-off, where improving tail performance can degrade performance on the overall dataset. Contrastive learning-based methods, such as Contrastive [makansi2021exposing] and FEND [wang2023fend], exhibit limited improvements in long-tail performance. For instance, on Argoverse 2 motion forecast dataset, contrastive learning [makansi2021exposing] achieves a minADE6 of 1.44 m and a minFDE6 of 2.74 m on top 10% samples. For the top 5% tail samples, the proposed CDKFormer achieves a 14.62% reduction on minFDE6, significantly surpassing FEND by a large margin. This advantage is even more pronounced on the inD dataset, where CDKFormer achieves a b-minFDE6 of 3.14 and 3.29 on top 10% and 5% tail samples, respectively. This validates the superior effectiveness of CDKFormer in handling diverse and challenging long-tail scenarios.
Figure 11: Long-tail performance comparison of QCNet and CDKFormer based on minFDE6 on Argoverse 2 motion forecast dataset.
Figure 12 further presents a detailed comparison of the error distributions for CDKFormer against long-tail learning baselines, evaluated on the top 10% of tail samples from the inD dataset. The violin plots show that the error distributions for CDKFormer are visibly shifted towards lower values across all four metrics: minADE6, minFDE6, b-minFDE6, and MR6. This qualitative observation is confirmed by the statistical significance analysis; a Mann-Whitney U test indicates that the improvements are statistically significant against nearly all baselines for the four metrics. This provides strong evidence of CDKFormer’s effectiveness in the rarest and most difficult scenarios.
(a) Distribution of minADE6
(b) Distribution of minFDE6
(c) Distribution of b-minFDE6
(d) Distribution of MR6
Figure 12: Performance comparison of long-tail learning methods on inD dataset. The violin plots show the error distributions for CDKFormer and four baseline long-tail learning methods, including balancing sampling, loss reweighting, Contrastive, and FEND, on the top 10% of tail samples of inD dataset. Performance is evaluated across four metrics: (a) minADE6, (b) minFDE6, (c) b-minFDE6, and (d) MR6. Statistical significance between CDKFormer and each baseline was determined using a Mann-Whitney U test. (*: $p<0.05$, **: $p<0.01$, ***: $p<0.001$, ns: not significant)
To provide a more stable measure of worst-case performance beyond single-sample maximums, we adopt conditional value at risk (CVaR), which quantifies the expected error in the tail of the distribution by averaging all error values that exceed a certain percentile threshold ($\alpha$). In this study, we conduct a granular analysis by applying CVaR to the minFDE6 metric at $\alpha$ levels from 90% to 99% in single-percentile increments. This allows us to evaluate the average model performance on progressively smaller slices of the worst-case scenarios, from the top 10% down to the top 1%. The results of the CVaR analysis are presented in Table 6, revealing CDKFormer’s superior ability to long-tail prediction. Our model achieves the lowest CVaR across the vast majority of the tail spectrum, from the 90th to the 97th percentile. For example, compared to the balanced sampling technique, CDKFormer achieves a 13.5% and 10.8% decrease in CVaR(minFDE6) on the worst 5% and 1% of tail samples, respectively. Interestingly, the loss reweighting method exhibits the strongest performance on the most extreme outliers (the 98th and 99th percentiles), while CDKFormer achieves the second-best performance. In general, CDKFormer provides a more consistent and substantial performance improvement across the broader set of difficult long-tail cases.
Table 6: CVaR for minFDE6 on the inD Dataset. The table shows the average minFDE6 for increasingly challenging subsets of the tail distribution, from the worst 10% (90th percentile) to the worst 1% (99th percentile) of cases.
CVaR at Percentile $\alpha$
90% 91% 92% 93% 94% 95% 96% 97% 98% 99%
QCNet[zhou2023query] 4.62 4.87 5.16 5.49 5.89 6.33 6.88 7.57 8.60 10.45
QCNet + Balanced Sampling 4.57 4.81 5.08 5.38 5.74 6.17 6.71 7.38 8.41 10.27
QCNet + Loss Reweighting 4.60 4.76 4.95 5.15 5.39 5.66 6.01 6.47 7.08 8.20
QCNet + Contrastive[makansi2021exposing] 4.62 4.87 5.14 5.45 5.84 6.29 6.88 7.63 8.80 11.24
QCNet + FEND[wang2023fend] 4.63 4.88 5.17 5.49 5.89 6.35 6.96 7.74 8.94 11.40
CDKFormer 3.88 4.07 4.33 4.59 4.89 5.34 5.80 6.44 7.53 9.17
5.4 Ablation Study
5.4.1 Ablation Study on Input Deviation Feature